# “partially massless” graviton, and photon fields in de Sitter spacetime under conformal symmetry

### v1

@@ IV.4 Remarks on related literature Takook
to:
[…]
Note that in Ref. Takook , the dS radius of curvature is set to $1$ ($R=1$). While this straightforward result does not, in itself, prove the subsequent computations incorrect, it raises concerns that the conclusions of Takook et al. could be significantly altered or simplified.
IV.5 The importance of global symmetry in dS physics
[…]
The global framework, as emphasized in this work, is far from being a mere abstract mathematical construct. It encapsulates critical aspects of dS physics, including the thermodynamic properties of the horizon, the dynamics of quantum fields, and the compatibility of quantum gravity with conformal symmetry. Specifically, the discussion in this paper centers on Dirac’s six-cone formalism, which provides a unified treatment of massless and “partially massless” fields in dS spacetime under conformal invariance. This formalism inherently depends on the global structure and cannot be meaningfully reduced to specific patches, underscoring the indispensable role of global symmetry in preserving the coherence of the physical framework.
Moreover, recent research by the authors EPL highlights the critical role of global symmetry in maintaining the coherence of dS spacetime. This study reveals that the matter-antimatter asymmetry, often regarded as a fundamental cosmological challenge, can manifest as an observer-dependent effect arising from the time orientation within local causal patches. Crucially, this asymmetry does not stem from an intrinsic property of dS spacetime but emerges due to the limitations of local perspectives. By leveraging the unique causal and analytic structure of the dS manifold, the study demonstrates that global symmetry considerations reconcile this apparent asymmetry without contradicting conventional mechanisms such as baryon number violation and CP violation. These findings underscore the indispensable role of a global, symmetry-based framework, not only in preserving the integrity of dS spacetime but also in addressing pivotal questions at the intersection of quantum field theory and cosmology.
@@ Acknowledgements
Appendix A Derivation and detailed calculations for Eq. (93)
@@
After performing a series of straightforward computations (utilizing Eqs. (70) and (III.1)) and rearranging the above result, it becomes evident that all terms involving $\boldsymbol{s_{2}}$ cancel out, leaving us with:
@@
Appendix B Derivation and detailed calculations for Eqs. (III.4) and (102)
Let us start with the traceless part, which satisfies Eq. (96), subject to the constraint (98). Following a similar approach as outlined in the vector case (see Appendix A), we define a ten-dimensional space $\mathfrak{E}$, which is spanned by the linear combinations of the following set of ten fundamental functions:
@@
After carrying out a series of straightforward calculations (utilizing Eqs. (70) and (III.1)) and reorganizing the result, it becomes clear that all terms involving $\boldsymbol{e_{7}}$ cancel out, leaving us with the following simple representation:
@@
Similar to the vector case (see Appendix A), due to the commutative nature of the operators $Q_{0}$ and $\left(Q_{0}-2\right)$, we can consider the following equation instead of starting with Eq. (96):
@@
After performing straightforward calculations (employing Eqs. (70) and (III.1)), we can simplify the result as shown below:

### v2

@@ IV.4 The importance of global symmetry in dS physics
The global framework, as emphasized in this work, is far from a purely formal or abstract construct. It encapsulates critical aspects of dS physics, including the thermodynamic properties of the horizon, the dynamics of quantum fields, and the compatibility of quantum gravity with conformal symmetry. Specifically, the discussion in this paper centers on Dirac’s six-cone formalism, which provides a unified treatment of massless and “partially massless” fields in dS spacetime under conformal invariance. This formalism inherently depends on the global structure and cannot be meaningfully reduced to specific patches. This underscores the indispensable role of global symmetry in preserving the coherence of the physical framework.
Moreover, recent research by the authors EPL highlights the crucial role of global symmetry in maintaining the internal consistency of dS spacetime. This study reveals that the matter-antimatter asymmetry, often regarded as a fundamental cosmological challenge, can manifest as an observer-dependent effect arising from the time orientation within local causal patches. This asymmetry does not stem from an intrinsic property of dS spacetime but emerges due to the limitations of local perspectives. By leveraging the unique causal and analytic structure of the dS manifold, the study demonstrates that global symmetry considerations reconcile this apparent asymmetry without contradicting established mechanisms such as baryon number violation and CP violation. This exemplifies how certain physical effects may appear differently—or even paradoxically—when analyzed solely within a local patch.
It is, however, essential to clarify that this emphasis on global symmetry does not diminish the value of the static patch approach. On the contrary, the two perspectives should be regarded as complementary. The global viewpoint provides a comprehensive framework for understanding symmetries and field dynamics on the entire manifold, while the static patch remains indispensable for exploring locally accessible phenomena, particularly those relevant to cosmological observations. Together, they form a multifaceted toolkit that enables a richer and more complete understanding of dS physics from different, yet compatible, vantage points.
@@ Acknowledgements
Appendix A Some useful relations
Here are some important relations crucial for verifying the equations presented in this article:
$\partial_{2}\cdot\theta\phi$ $=$ $-R^{-2}D_{1}\phi\,,$ (1)
$Q_{1}D_{1}\phi$ $=$ $D_{1}Q_{0}\phi\,,$ (2)
$\partial_{2}\cdot D_{2}D_{1}\phi$ $=$ $-\left(Q_{1}+6\right)D_{1}\phi$ (3)
$=$ $-D_{1}\left(Q_{0}+6\right)\phi\,,$
where $\phi$ is a scalar field on dS spacetime.
Moreover, we have the following relations:
$\left(Q_{0}-2\right)x_{a}=$ $\,x_{a}\left(Q_{0}-2\right)-4x_{a}-2D_{1a}\,,$ (4)
$\left(Q_{0}-2\right)D_{1a}=$ $\,D_{1a}\left(Q_{0}-2\right)+6D_{1a}$
$\,+2\left(Q_{0}+4\right)x_{a}\,,$ (5)
$Q_{0}\left(Q_{0}-2\right)x_{a}=$ $\,x_{a}Q_{0}\left(Q_{0}-2\right)$
$\,-4(3x_{a}+D_{1a})\left(Q_{0}-2\right)\,,$ (6)
$D_{1a}D_{1b}=$ $\,D_{1b}D_{1a}+x_{b}D_{1a}-x_{a}D_{1b}\,.$ (8)
Appendix B Link to intrinsic coordinates
$h_{\mu\nu}(X)=x^{a}_{\,\,,\,\mu}x^{b}_{\,\,,\,\nu}\;{\cal{K}}_{ab}\big{(}x(X)\big{)}\,,$ (9)
The following transformation applies to the covariant derivatives:
$=$ ¯$\bar{\partial}^{2}\;\phi=-R^{-2}Q_{0}\,,$
where $\phi$ denotes a dS scalar field.
Appendix C Derivation and detailed calculations for Eq. (80)
@@ Appendix B Link to intrinsic coordinates
After performing a series of straightforward computations (utilizing Eqs. (4) and (A)) and rearranging the above result, it becomes evident that all terms involving $\boldsymbol{s_{2}}$ cancel out, leaving us with:
@@
Appendix D Derivation and detailed calculations for Eqs. (III.4) and (90)
Let us start with the traceless part, which satisfies Eq. (83), subject to the constraint (85). Following a similar approach as outlined in the vector case (see Appendix C), we define a ten-dimensional space $\mathfrak{E}$, which is spanned by the linear combinations of the following set of ten fundamental functions:
@@
After carrying out a series of straightforward calculations (utilizing Eqs. (4) and (A)) and reorganizing the result, it becomes clear that all terms involving $\boldsymbol{e_{7}}$ cancel out, leaving us with the following simple representation:
@@
Similar to the vector case (see Appendix C), due to the commutative nature of the operators $Q_{0}$ and $\left(Q_{0}-2\right)$, we can consider the following equation instead of starting with Eq. (83):
@@
After performing straightforward calculations (employing Eqs. (4) and (A)), we can simplify the result as shown below: